Is (9, 1) a solution to this system of equations?
x + 3y = 12
18x + 16y = 9
yes
no
Submit


Answer by t0hierry(194)

no
it needs to satisfy both equations
18x + 54y = 216
18x + 16y = 9
38y = 207
y = 207/38